# StreamingThinker: Large Language Models Can Think While Reading

### v1

@@ Appendix B Streaming Thinking Paradigm
Human reasoning during conversation or reading naturally unfolds in a streaming manner—people engage in comprehension and inference as information arrives, without waiting for the complete context. Inspired by this cognitive process, we proposed in the main text a Streaming Thinking Paradigm that enables large language models to reason concurrently with incremental input. In this appendix, we further elaborate on this paradigm by providing detailed explanations and illustrative examples that clarify its operational design and the ordering of context and question.
At each step, the model incrementally processes the incoming sentence, focusing on progressive comprehension: (1) understanding and summarizing key information, (2) explaining ambiguities and reorganizing semantic relations, (3) extending logical implications, and (4) skipping thinking step when the content is irrelevant to the question. After completing incrementally reading and reasoning, the model may (1) directly produce an answer, representing the shallowest depth; (2) further integrate global information to achieve deeper comprehension; or (3) perform reflective reasoning on top of global integration to obtain the most reliable reasoning outcome.
B.1 Examples of Paradigm Design
@@
Given a multi-sentence word problem, the model reads the context sentence by sentence, identifying rate/quantity information, skipping irrelevant text, and maintaining interpretable intermediate results (e.g., partial counts). The same pipeline supports three depths of answers: D1 a direct computation from the current notes, D2 a globally consolidated solution that aggregates all intervals, and D3 a reflective pass that re-checks each arithmetic step to ensure reliability.
@@
B.2 Examples of the Ordering of Context and Question
@@ C.1 Does Incremental Sentence Input Work?
A straightforward idea is to feed sentences into LLMs sequentially and expect sentence-by-sentence reasoning based on the accumulated input. However, our experiments show that when given only a single sentence without streaming fine-tuning, the original reasoning model tends to overthink (Sui et al., 2025). In other words, although this sentence-level data generation method appears to align better with the streaming property, the vanilla model lacks such generative capability and exhibits reduced controllability in CoT generation.
@@ C.4 Quality Evaluation of Streaming CoT
where and denote the embedding vectors of reasoning sentence and input sentence , respectively. We employ SentenceBERT (Reimers & Gurevych, 2019) to obtain embeddings. A higher score reflects that the reasoning faithfully preserves both the order and semantic content of the input stream. In cases where multiple reasoning sentences correspond to one input sentence, we aggregate them as a single segment.
@@ Appendix D Dataset Details
We construct math reasoning data from GSM-Symbolic (Mirzadeh et al., 2025), MetaMathQA (Yu et al., 2024), and TuLu-personas-math-grade (Lambert et al., 2024).
GSM-Symbolic is a symbolic variant of GSM8K that reformulates arithmetic word problems into semi-structured symbolic expressions, explicitly preserving quantities, operators, and logical relations. It consists of two subsets, GSM-Symbolic-P1 (5K samples) and GSM-Symbolic-P2 (2.5K samples), where P2 includes longer and more complex problems. We randomly sample 4K and 2K instances from P1 and P2 for training, respectively, while the remaining data are used for generating training instances.
@@
For logical reasoning, we evaluate on ProofWriter (Tafjord et al., 2020) and LogicNLI (Tian et al., 2021). ProofWriter is a benchmark designed for evaluating multi-step logical reasoning and theorem proving capabilities in natural language. It is derived from the EntailmentBank corpus, where each example consists of a hypothesis and a set of supporting facts expressed in natural language. The task requires the model to infer whether the hypothesis is entailed, contradicted, or neutral given the supporting facts, while optionally generating an explicit reasoning chain that connects the premises to the conclusion. LogicNLI is a natural language inference (NLI) benchmark designed to evaluate the logical reasoning ability of language models beyond surface-level semantics. Each example in LogicNLI consists of a premise and a hypothesis pair, annotated with one of three logical relations: entailment, contradiction, or neutral. Unlike conventional NLI datasets that focus primarily on lexical or syntactic cues, LogicNLI emphasizes reasoning over formal logic structures such as conjunction, disjunction, negation, quantifiers, and implication.
@@
For context-based QA reasoning task, we evaluate StreamingThinker on PubMedQA (Jin et al., 2019) and HotpotQA (Yang et al., 2018) datasets.
@@ E.1 Training Details
For training, we set the batch size to 16 and employ the AdamW optimizer with a learning rate of 1e-5, a cosine decay schedule, and a 300-step linear warm-up. We additionally enable activation checkpointing to mitigate memory overhead from maintaining parallel caches during training. These choices align the training setup with prior LLM pretraining while introducing only the modifications necessary for streaming compatibility.
@@ E.2 Decoding Strategy
For generation, we follow the sampling parameters of Qwen3. We adopt nucleus sampling () with temperature and . This balances fluency and diversity while maintaining consistency across sequentially produced reasoning sentences. Importantly, because input and reasoning caches are maintained independently, decoding proceeds with lower latency: new inputs can be consumed without flushing or rewriting the reasoning cache, and reasoning updates can be emitted as soon as the corresponding input sentence completes. This strategy preserves alignment with streaming training while ensuring responsiveness in real-time scenarios. The streaming thinking decoding process is shown in Algorithm 1.
@@ E.3 Relation between Parallel KV Caches and Prefill-Decode Separation
While our parallel KV caches mechanism may formally resemble techniques for Prefill-Decode separation (Zhong et al., 2024; Qin et al., 2024), the foundational motivations and operational goals behind the two approaches are fundamentally distinct.

### v2

@@ Appendix B Streaming Thinking Paradigm
B.1 Cognitive Foundations of Streaming Thinking
Human reasoning during reading naturally unfolds in a streaming manner—people engage in comprehension and inference as information arrives, without waiting for the complete context.
Our proposed StreamingThinker architecture draws structural inspiration from established models of human discourse comprehension, specifically the Construction-Integration (CI) model proposed by Kintsch (kintsch1988role). The CI model posits that comprehension occurs in two alternating cycles:
The Construction Phase: A bottom-up process where linguistic input triggers the activation of concepts and propositions based on local context. In this stage, the cognitive system acts as a high-bandwidth information buffer, prioritizing the establishment of local coherence (e.g., resolving immediate pronouns or connecting adjacent clauses) over global consistency. Crucially, this phase is non-selective: it allows multiple, potentially conflicting interpretations to co-exist in a temporary cognitive state. This ensures that no critical information is prematurely discarded before the full context is available
The Integration Phase: Once the initial propositional network is constructed, the cognitive system iteratively updates node activations based on their connectivity and connection strengths. Through this process, the system resolves local ambiguities (e.g., polysemy) and filters out contextually inappropriate inferences. The result is a refined macrostructure where only the information strictly consistent with the global context remains active, ensuring a unified and logical understanding of the discourse.
Inspired by this cognitive process, we proposed in the main text a Streaming Thinking Paradigm that enables large language models to reason concurrently with incremental input. Our streaming phase corresponds to the construction phase, where the model performs ”shallow reasoning” (e.g., entity tracking, intermediate calculation) to process incoming chunks and maintain local coherence (graesser1994constructing). Our final reasoning phase corresponds to the integration phase, where the model utilizes the fully accumulated context (KV cache) to synthesize a globally consistent answer. This theoretical alignment explains why our model avoids the ”hallucination” pitfalls of premature guessing: like a human reader, it defers the final integration of complex causal chains until the necessary global information is available.
In this appendix, we further elaborate on this paradigm by providing detailed explanations and illustrative examples that clarify its operational design and the ordering of context and question.
At each step, the model incrementally processes the incoming sentence, focusing on shallow and progressive comprehension, which aligns the construction phase in human comprehension. Specifically, we designed distinct categories of tasks for the model during local construction, serving as cognitive scaffolds for establishing local coherence. These tasks (e.g., intermediate calculation, entity tracking) compel the model to explicitly process and encode immediate details, transforming raw input into structured representations without prematurely committing to a global conclusion. (1) understanding and summarizing key information, (2) explaining ambiguities and reorganizing semantic relations, (3) extending logical implications, and (4) skipping thinking step when the content is irrelevant to the question.
After completing incrementally reading and reasoning, the model shifts its focus from local coherence to global interpretation. It leverages the full context now available in its memory (KV cache) to perform the deep, unconstrained reasoning that was deliberately deferred during the streaming phase. Specifically, the model may: (1) directly produce an answer, representing the shallowest depth; (2) further integrate global information to achieve deeper comprehension; or (3) perform reflective reasoning on top of global integration to obtain the most reliable reasoning outcome.
B.2 Examples of Paradigm Design
@@
Given a multi-conditions problem, the model reads the context sentence by sentence, identifying rate/quantity information, skipping irrelevant text, and maintaining interpretable intermediate results (e.g., partial counts). The pipeline supports three reasoning depths: D1 a direct computation from the current notes, D2 a globally consolidated solution that aggregates all intervals, and D3 a reflective pass that re-checks each arithmetic step to ensure reliability.
@@
B.3 Examples of the Ordering of Context and Question
@@ C.1 Does Incremental Sentence Input Work?
A straightforward idea is to feed sentences into LLMs sequentially and expect sentence-by-sentence reasoning based on the accumulated input. However, our experiments show that when given only a single sentence without streaming fine-tuning, the original reasoning model tends to overthink (sui2025stop). In other words, although this sentence-level data generation method appears to align better with the streaming property, the vanilla model lacks such generative capability and exhibits reduced controllability in CoT generation.
@@ C.4 Quality Evaluation of Streaming CoT
where and denote the embedding vectors of reasoning sentence and input sentence , respectively. We employ SentenceBERT (reimers2019sentence) to obtain embeddings. A higher score reflects that the reasoning faithfully preserves both the order and semantic content of the input stream. In cases where multiple reasoning sentences correspond to one input sentence, we aggregate them as a single segment.
@@ Appendix D Dataset Details
We construct math reasoning data from GSM-Symbolic (mirzadehgsm), MetaMathQA (yumetamath), and TuLu-personas-math-grade (lambert2024tulu).
GSM-Symbolic is a symbolic variant of GSM8K that reformulates arithmetic word problems into semi-structured symbolic expressions, explicitly preserving quantities, operators, and logical relations. It consists of two subsets, GSM-Symbolic-P1 (5K samples) and GSM-Symbolic-P2 (2.5K samples), where P2 includes longer and more complex problems. We randomly sample 4K and 2K instances from P1 and P2 for training, respectively, while the remaining data are used for evaluation.
@@
For logical reasoning, we evaluate on ProofWriter (tafjord2020proofwriter) and LogicNLI (tian2021diagnosing). ProofWriter is a benchmark designed for evaluating multi-step logical reasoning and theorem proving capabilities in natural language. It is derived from the EntailmentBank corpus, where each example consists of a hypothesis and a set of supporting facts expressed in natural language. The task requires the model to infer whether the hypothesis is entailed, contradicted, or neutral given the supporting facts, while optionally generating an explicit reasoning chain that connects the premises to the conclusion. LogicNLI is a natural language inference (NLI) benchmark designed to evaluate the logical reasoning ability of language models beyond surface-level semantics. Each example in LogicNLI consists of a premise and a hypothesis pair, annotated with one of three logical relations: entailment, contradiction, or neutral. Unlike conventional NLI datasets that focus primarily on lexical or syntactic cues, LogicNLI emphasizes reasoning over formal logic structures such as conjunction, disjunction, negation, quantifiers, and implication.
@@
For context-based QA reasoning task, we evaluate StreamingThinker on PubMedQA (jin2019pubmedqa) and HotpotQA (yang2018hotpotqa) datasets.
@@ E.1 Training Details
For training, we set the batch size to 16 and employ the AdamW optimizer with a learning rate of 1e-5. We additionally enable activation checkpointing to mitigate memory overhead from maintaining parallel caches during training. These choices align the training setup with prior LLM pretraining while introducing only the modifications necessary for streaming compatibility.
@@ E.2 Decoding Strategy
For generation, we follow the sampling parameters of Qwen3. This balances fluency and diversity while maintaining consistency across sequentially produced reasoning sentences. Importantly, because input and reasoning caches are maintained independently, decoding proceeds with lower latency: new inputs can be consumed without flushing or rewriting the reasoning cache, and reasoning updates can be emitted as soon as the corresponding input sentence completes. This strategy preserves alignment with streaming training while ensuring responsiveness in real-time scenarios. The streaming thinking decoding process is shown in Algorithm 1.
@@ E.3 Relation between Parallel KV Caches and Prefill-Decode Separation
While our parallel KV caches mechanism may formally resemble techniques for Prefill-Decode separation (zhong2024distserve; qin2024mooncake), the foundational motivations and operational goals behind the two approaches are fundamentally distinct.